Isaac applies a force of 40N at an angle of 60°up from the horizontal to a wooden rod using a spring scale.If he generatesa torque of 73 m.N,how long was the rod?

1
Expert's answer
2021-11-18T10:16:32-0500

"F = 40N"

"\alpha = 60\degree"

"\tau=73N*m"

"\tau = F*r"

"\tau=|r|*|F|*\sin \alpha"

"r= \frac{\tau}{F*\sin \alpha}=\frac{73}{40*\sin 60\degree}\approx2.11m"


"\text{Answer:}2.11m"